OpenStudy (anonymous):

solve this it is really hard can nybody help me I WILL GIVE A MEADEL AND FAN. (9^62773 + 2)^83721= 83721 x log (9^62773 + 2)

OpenStudy (ahsome):

\[(9^62773 + 2)^{83721}= 83721\times log(9^{62773} + 2) \] That it?

OpenStudy (anonymous):

YEP THATS IT IAM IN COLLGE I REALY NEED HELP

OpenStudy (ahsome):

Boy, you need to use a calculator for this one

OpenStudy (ahsome):

\[(9^{62773} + 2)^{83721}= 83721\times log (9^{62773} + 2) \]

OpenStudy (anonymous):

IS THIS RIGHT can you check this log (9^62773 + 2) = log (9^62773) +log ( 2) log (9^62773 + 2) = 62773 log (9) + log ( 2) log (9^62773 + 2) = 62773 (0.954242509) + ( 0.301029996) log (9^62773 + 2) = 59900.966 So: 83721 x log (9^62773 + 2)= (83721)(59900.966) 83721 x log (9^62773 + 2) =5.01496877 × 10^9 Digital sum is 5 + 0+ 1+4+9+6+8+7+7= 47 4+7=11 1+1=2

OpenStudy (anonymous):

ive been working on this all night

OpenStudy (ahsome):

Hmm...

OpenStudy (ahsome):

Its on Yahoo! Answers, so I guess so?

OpenStudy (anonymous):

i did not det that from yahoo sorry my collge math teacher gave me this problem because i am in advance math so i gusse he thought i could do this for some reason

OpenStudy (anonymous):

and did u know my cousin play for noterdame he is number 5 in the picture

OpenStudy (ahsome):

Really? NIce